# A Novel, Steerable, Low-Energy Proton Source for Detector Characterization

### v1

@@ Abstract
We report on the conversion of the Manitoba II mass spectrometer into a versatile proton beam facility. This infrastructure is adaptable to any detector-under-test (DUT), and has proven itself effective with the characterization of silicon detectors used in subatomic beyond-the-Standard-Model (BSM) searches. This work discusses the modifications and subsequent tests to confirm the beam specifications met the demands of the aforementioned detectors.
@@ 1 Introduction
Low energy, high precision neutron and neutrino experiments provide a rich arena to probe BSM physics [1]. Many of these experiments such as the Nab, pNAB, and Katrin experiments use silicon-based detector technologies due to their high energy resolution, fast timing, insensitivity to magnetic fields, and their position sensitivity [2, 3, 4, 5, 6]. The University of Manitoba’s high-precision mass spectrometer (“Manitoba II”) was adapted to provide a unique test station for effectively any detector-under-test (DUT). Specifically, it was commissioned as a proton source to fill the testing requirements for the Neutron-ab (Nab) experiment, where a, and b are the parameters being measured.
The Nab detectors have a $117\text{\,}\mathrm{mm}$ diameter active area divided into 127 hexagonal pixels (each with $\sim$$10\text{\,}\mathrm{m}\mathrm{m}$ height). The proton beam was required to traverse the entire face of the detector, while maintaining a small “spot size”, less than $6.5\text{\,}{\mathrm{mm}}^{2}$ area to remain within the envelope of the pixelation, enabling single pixel study without significant charge sharing between the pixels. In what follows, we break down the stages within the proton source (Section 2), going through each segment in turn, and then discuss the studies performed to understand the beam parameters under typical operation (see Section 3).
@@ 2 Components of the Proton Source
The Manitoba II mass spectrometer was a double-focusing (velocity and angle) high precision instrument built in 1967 for precise mass measurements of positively charged ions [7]. After serving the atomic physics community for over forty years, it gained a second lease on life through the adaptation to a steerable low-energy “pencil beam” proton source for detector characterization [8]. Figure 1 shows a schematic of the proton source facility after modifications. A gas discharge source is used to produce an ion beam, replacing the old tantalum oven-ized source (see 2.1). The diffusion pumps were removed from the vacuum system to reduce the risk of backward contamination to sensitive detector instrumentation in favor of more conventional turbo-molecular pumps, backed by diaphragm roughing pumps. The ion beam then passes through the unmodified Electro-static Analyzer (ESA) (see 2.2) and Magneto-static Analyzer (MSA) (see 2.3). An electrostatic steerer was installed after the MSA to deflect the proton beam over the face of a detector in the Detector Region. (see 2.4). The Detector Region includes a vacuum cross with a long cylindrical vacuum vessel downstream which accepts the $1\text{\,}\mathrm{m}$ long Nab silicon detector, amplifier, and readout assembly [9, 2].
@@ 2.1 Penning Ion Generator
A Penning ion generator (PIG) source was installed to produce protons for use in detector characterization experiments (see Figure 2) [8, 10]. A hydrogen-argon gas mixture is introduced into the high vacuum discharge region where the rate is controlled by a manual needle valve. The argon gas in the mixture functions as a buffer gas, limiting the mobility of free charge in the discharge vacuum region. This has the effect of regulating the source discharge current and ultimately the stability of the discharge, resulting in a continuous discharge. The relative concentrations of hydrogen and argon gases in the gas mixture were empirically determined by introducing increasing quantities of hydrogen gas until the source arc current was stable at $\sim$ 1-$3\text{\,}\mathrm{mA}$.
@@
In the discharge region, once an ion reaches the anode/cathode it no longer participates in ionizing atoms/molecules. A high critical-temperature permanent magnet (42% Neodymium) is used to increase the time spent by ions ionizing other particles in the discharge region. Charged particles follow helical trajectories in the electromagnetic field and may even be circulated across the discharge gap multiple times before extinction depending on the ion’s initial conditions and the space-charge effects of the plasma. Trajectories with large path lengths relative to their extinction time will have more time to ionize particles thereby increasing ion production in the discharge region. Ions near the extraction plate are accelerated across the $30\text{\,}\mathrm{kV}$ acceleration gap toward the extraction tip before moving to the ESA for energy analysis. The amount of energy imparted upon the accelerated ions depends on the type of ion. Since protons are singly charged ions they gain $30\text{\,}\mathrm{k}\mathrm{e}\mathrm{V}$ of energy when accelerated through the gap, although we note that ionic $H^{+}_{2}$ and and $H_{3}^{+}$ ions are also created (see Figure 5).
@@ 2.2 Electro-static Analyzer
The electro-static analyzer (ESA) consists of two concentric, arced plates to analyze ions according to their kinetic energy (see Figure 1). The ions travel in an arced path of radius $1\text{\,}\mathrm{m}$ due to the radial electric field produced between the plates. This field is generated by applying a constant deflection potential $V_{d}\equiv V_{b}-V_{a}$ across the $2\text{\,}\mathrm{c}\mathrm{m}$ electrode gap, where the outer electrode is held at $V_{b}$ and the inner electrode is held at $V_{a}$, while requiring that $V_{b}=-V_{a}$ [11]. Ions that possess the selected energy travel along the central circular trajectory of the ESA because the ion experiences a centripetal force $\vec{F}_{C}$ that is equal in magnitude and opposite in direction to the force of the radial electric field $\vec{F}_{E}$; thus there is no net force on ions traveling the central trajectory. Ions possessing energies other than the selected energy will experience a radial force which causes a them to travel a trajectory deviating from that of the central trajectory. For $30\text{\,}\mathrm{k}\mathrm{e}\mathrm{V}$ ions, $V_{d}=$-1200\text{\,}\mathrm{V}$\$. A collimating slit at the exit of the ESA ensures only ions traveling the central circular path leave the ESA which effectively selects ions by their initial kinetic energy.
@@ 2.4 Proton Steerer
The electro-static steerer is located in-between the MSA and the detector region. As shown in Figure 3, the steerer consists of a set of 4 independent rectangular copper electrodes and is placed in a 10” long 8” Conflat nipple. Application of static potentials of up to $\pm 2\text{\,}\mathrm{kV}$ can be individually applied to the parallel plates of the steerer. The dimensions and voltages of the steerer were determined by simulating proton trajectories through different electrode geometries using the SIMION ion optics simulation software such that protons would be deflected up to $75\text{\,}\mathrm{mm}$ radially at the Nab silicon detector location $0.48\text{\,}\mathrm{m}$ from exit of the electron-static steerer [12].
@@ 3.1 Proton Energy Resolution
Figure 5: A Micro-channel plate (MCP) detector measures counts per minute with pencil beams of various hydrogen species passing through both analysis sectors while varying the magnetic field in the MSA. The solid vertical lines are calculated using a $27.245\text{\,}\mathrm{kV}$ acceleration voltage and Equation 1. This acceleration voltage can be varied within a range of stability — for this test $27.245\text{\,}\mathrm{kV}$ was the highest stable voltage, however improvements were made to achieve up to $35\text{\,}\mathrm{kV}$. Also note that the count rate is highly dependent on the PIG configuration, and the noise in the detector used [8].
@@
Thus the error in our proton energy is $\sim$$300\text{\,}\mathrm{e}\mathrm{V}$ FWHM, which is generally much lower than the energy resolution of the Nab silicon detectors [9].
Figure 6: Proton counts as a function of magnetic field strength in the MSA to resolve the $H^{+}$ peak. The red vertical lines represent $2\text{\,}\mathrm{G}$ FWHM. This data translates to a proton energy resolution of $300\text{\,}\mathrm{e}\mathrm{V}$ via Equation 4. This is much lower than the energy resolution of the Nab silicon detectors, allowing a precise confirmation of the Nab detector performance [9].
@@ 4 Conclusion
The Manitoba II spectrometer was modified with a $30\text{\,}\mathrm{k}\mathrm{e}\mathrm{V}$ accelerating potential penning trap ion source and a electro static steerer with the goal of characterizing the 117 mm diameter, 127 pixel Nab silicon detectors. The spectrometer easily isolates the $30\text{\,}\mathrm{k}\mathrm{e}\mathrm{V}$ protons generated from the penning ion generator with an energy resolution of $300\text{\,}\mathrm{e}\mathrm{V}$ FHWM. During the testing of the Nab detectors, We ran at $10\text{\,}\mathrm{p}\mathrm{r}\mathrm{o}\mathrm{t}\mathrm{o}\mathrm{n}\mathrm{s}\mathrm{/}\mathrm{s}$, to reduce electrical noise in the data acquisition system. To evaluate the proton spot size, a large diameter phosphor screen was employed where most spot sizes in the area of interest were $1\text{\,}\mathrm{m}\mathrm{m}^{2}$ area with a maximum area of $5\text{\,}{\mathrm{mm}}^{2}$. In addition the spot size was evaluated by scanning over a pixel on the Nab silicon detector which determined a spot size diameter of $3.1\text{\,}{\mathrm{mm}}^{2}$. This demonstrates that the Manitoba II proton source met the testing requirements for the Nab silicon detectors and is a competent testing apparatus for calibration and characterization of low-energy physics particle detectors.

### v2

@@ Abstract
We report on the conversion of the Manitoba II mass spectrometer into a versatile low-energy proton beam facility. This infrastructure is adaptable to any detector-under-test (DUT), and has proven itself effective with the characterization of silicon detectors used in subatomic beyond-the-Standard-Model (BSM) searches, namely the Nab experiment. A pencil beam of mono-energetic protons can be produced in a range from $25\text{\,}\mathrm{k}\mathrm{e}\mathrm{V}$ to $35\text{\,}\mathrm{k}\mathrm{e}\mathrm{V}$, achieving a beam current of $\sim$$1\text{\times}{10}^{-18}\text{\,}\mathrm{A}$. Electrostatic steering plates were constructed to direct the Gaussian-profile proton beam over a $117\text{\,}\mathrm{mm}$ diameter area-of-interest with full-width at half-maxima (FWHM) ranging from $0.6\text{\,}\mathrm{mm}$ to $1.26\text{\,}\mathrm{mm}$. This work discusses the modifications and subsequent tests to confirm the beam specifications met the demands of the aforementioned detectors.
@@ 1 Introduction
Low energy, high precision neutron and neutrino experiments provide a rich arena to probe BSM physics [FSNNwhitepaper2023]. Many of these experiments such as the Nab, pNAB, and Katrin experiments use silicon-based detector technologies due to their high energy resolution, fast timing, insensitivity to magnetic fields, and their position sensitivity [Fry_2019, Baessler_2024, PSTP2024PNab, mertensNovelDetectorSystem2019, urbanCharacterizationMeasurementsTRISTAN2022]. The University of Manitoba’s high-precision mass spectrometer (“Manitoba II”) was adapted to provide a unique test station for effectively any detector-under-test (DUT). Specifically, it was commissioned as a proton source to fill the testing requirements for the Neutron-ab (Nab) experiment, where $a$, and $b$ are the parameters being measured.
The Nab detectors have a $117\text{\,}\mathrm{mm}$ diameter active area divided into 127 hexagonal pixels (each with $\sim$$10\text{\,}\mathrm{m}\mathrm{m}$ height). The proton beam was required to traverse the entire face of the detector, while maintaining a small “spot size”, less than $6.5\text{\,}{\mathrm{mm}}^{2}$ area to remain within the envelope of the pixelation, enabling single pixel study without significant charge sharing between the pixels. The performance and analysis of the Nab silicon detectors using the facility is reported here [nabsilicon_uofm], while this work focuses on the development and performance of the proton source used for these studies. In what follows, we break down the stages within the proton source (Section 2), going through each segment in turn, and then discuss the studies performed to understand the beam parameters under typical operation (see Section 3).
@@ 2 Components of the Proton Source
The Manitoba II mass spectrometer was a double-focusing (velocity and angle) high precision instrument built in 1967 for precise mass measurements of positively charged ions [Barber1971]. After serving the atomic physics community for over forty years, it gained a second lease on life through the adaptation to a steerable low-energy “pencil beam” proton source for detector characterization [Harrison2013]. Figure 1 shows a schematic of the proton source facility after modifications. A gas discharge source is used to produce an ion beam, replacing the old tantalum oven-ized source (see 2.1). The diffusion pumps were removed from the vacuum system to reduce the risk of backward contamination to sensitive detector instrumentation in favor of more conventional turbo-molecular pumps, backed by diaphragm roughing pumps. The ion beam then passes through the unmodified Electro-static Analyzer (ESA) (see 2.2) and Magneto-static Analyzer (MSA) (see 2.3). An electrostatic steerer was installed after the MSA to deflect the proton beam over the face of a detector in the Detector Region. (see 2.4). The Detector Region includes a vacuum cross with a long cylindrical vacuum vessel downstream which accepts the $1\text{\,}\mathrm{m}$ long Nab silicon detector, amplifier, and readout assembly [Broussard2017, Fry_2019].
@@ 2.1 Penning Ion Generator
A Penning ion generator (PIG) source was installed to produce protons for use in detector characterization experiments (see Figure 2) [Harrison2013, NickThesis]. A hydrogen-argon gas mixture is introduced into the high vacuum discharge region where the rate is controlled by a manual needle valve. The argon gas in the mixture functions as a buffer gas, limiting the mobility of free charge in the discharge vacuum region. This has the effect of regulating the source discharge current and ultimately the stability of the discharge, resulting in a continuous discharge. The relative concentrations of hydrogen and argon gases in the gas mixture were empirically determined by introducing increasing quantities of hydrogen gas until the source arc current was stable at $\sim$ 1-$3\text{\,}\mathrm{mA}$.
@@
In the discharge region, once an ion reaches the anode/cathode it no longer participates in the ionization process. A high critical-temperature permanent magnet (42% Neodymium) is used to increase the time spent by ions ionizing other particles in the discharge region. Charged particles follow helical trajectories in the electromagnetic field and may even be circulated across the discharge gap multiple times before extinction depending on the ion’s initial conditions and the space-charge effects of the plasma. Trajectories with large path lengths relative to their extinction time will have more time to ionize particles thereby increasing ion production in the discharge region. Ions near the extraction plate are accelerated across the $30\text{\,}\mathrm{kV}$ acceleration gap toward the extraction tip before moving to the ESA for energy analysis. The amount of energy imparted upon the accelerated ions depends on the type of ion. Since protons are singly charged ions they gain $30\text{\,}\mathrm{k}\mathrm{e}\mathrm{V}$ of energy when accelerated through the gap, although we note that ionic $H^{+}_{2}$ and and $H_{3}^{+}$ ions are also created (see Figure 5).
@@ 2.2 Electro-static Analyzer
The electro-static analyzer (ESA) consists of two concentric, arced plates to analyze ions according to their kinetic energy (see Figure 1). The ions travel in an arced path of radius $1\text{\,}\mathrm{m}$ due to the radial electric field produced between the plates. This field is generated by applying a constant deflection potential $V_{d}\equiv V_{b}-V_{a}$ across the $2\text{\,}\mathrm{c}\mathrm{m}$ electrode gap, where the outer electrode is held at $V_{b}$ and the inner electrode is held at $V_{a}$, while requiring that $V_{b}=-V_{a}$ [Liebl2008]. Ions that possess the selected energy travel along the central circular trajectory of the ESA because the ion experiences a centripetal force $\vec{F}_{C}$ that is equal in magnitude and opposite in direction to the force of the radial electric field $\vec{F}_{E}$; thus there is no net force on ions traveling the central trajectory. Ions possessing energies other than the selected energy will experience a radial force which causes a them to travel a trajectory deviating from that of the central trajectory. For $30\text{\,}\mathrm{k}\mathrm{e}\mathrm{V}$ ions, $V_{d}=$-1200\text{\,}\mathrm{V}$\$. A collimating slit at the exit of the ESA ensures only ions traveling the central circular path leave the ESA which effectively selects ions by their initial kinetic energy.
@@ 2.4 Proton Steerer
The electro-static steerer is located in-between the MSA and the detector region. As shown in Figure 3, the steerer consists of a set of 4 independent rectangular copper electrodes and is placed in a 10” long 8” Conflat nipple. Application of static potentials of up to $\pm 2\text{\,}\mathrm{kV}$ can be individually applied to the parallel plates of the steerer. The dimensions and voltages of the steerer were determined by simulating proton trajectories through different electrode geometries using the SIMION ion optics simulation software such that protons would be deflected up to $75\text{\,}\mathrm{mm}$ radially at the Nab silicon detector location $0.48\text{\,}\mathrm{m}$ from exit of the electron-static steerer [simion].
@@
2.5 Beam Characteristics
The proton beam can be tuned to the desired energy by adjusting the acceleration potential in the PIG. For studies of the Nab detectors, $25\text{\,}\mathrm{k}\mathrm{e}\mathrm{V}$, $30\text{\,}\mathrm{k}\mathrm{e}\mathrm{V}$, and $35\text{\,}\mathrm{k}\mathrm{e}\mathrm{V}$ were selected. The beam has no time structure, and instead is a constant stream of protons, with a beam current on the order of $1\text{\times}{10}^{-18}\text{\,}\mathrm{A}$. As will be discussed below, the physical extent of the beam is quite small, less than $5\text{\,}\mathrm{mm}$ in diameter, meaning that beam profile was assumed to be Gaussian.
@@ 3.1 Proton Energy Resolution
Figure 5: A Micro-channel plate (MCP) detector measures counts per minute with pencil beams of various hydrogen species passing through both analysis sectors while varying the magnetic field in the MSA. The solid vertical lines are calculated using a $27.245\text{\,}\mathrm{kV}$ acceleration voltage and Equation 1. This acceleration voltage can be varied within a range of stability — for this test $27.245\text{\,}\mathrm{kV}$ was the highest stable voltage, however improvements were made to achieve up to $35\text{\,}\mathrm{kV}$. Also note that the count rate is highly dependent on the PIG configuration, and the noise in the detector used [Harrison2013].
@@
Thus the error in our proton energy is $\sim$$300\text{\,}\mathrm{e}\mathrm{V}$ FWHM, which is generally much lower than the energy resolution of the Nab silicon detectors [Broussard2017].
Figure 6: Proton counts as a function of magnetic field strength in the MSA to resolve the $H^{+}$ peak. The red vertical lines represent $2\text{\,}\mathrm{G}$ FWHM. This data translates to a proton energy resolution of $300\text{\,}\mathrm{e}\mathrm{V}$ via Equation 4. This is much lower than the energy resolution of the Nab silicon detectors, allowing a precise confirmation of the Nab detector performance [Broussard2017].
@@ 4 Conclusion
The Manitoba II spectrometer was modified with a $30\text{\,}\mathrm{k}\mathrm{e}\mathrm{V}$ accelerating potential penning trap ion source and a electro static steerer with the goal of characterizing the 117 mm diameter, 127 pixel Nab silicon detectors. The spectrometer easily isolates the $30\text{\,}\mathrm{k}\mathrm{e}\mathrm{V}$ protons generated from the penning ion generator with an energy resolution of $300\text{\,}\mathrm{e}\mathrm{V}$ FHWM. During the testing of the Nab detectors, we ran at $10\text{\,}\mathrm{p}\mathrm{r}\mathrm{o}\mathrm{t}\mathrm{o}\mathrm{n}\mathrm{s}\mathrm{/}\mathrm{s}$, to reduce electrical noise in the data acquisition system. To evaluate the proton spot size, a large diameter phosphor screen was employed where most spot sizes in the area of interest were $1\text{\,}\mathrm{m}\mathrm{m}^{2}$ area with a maximum area of $5\text{\,}{\mathrm{mm}}^{2}$. In addition the spot size was evaluated by scanning over a pixel on the Nab silicon detector which determined a spot size diameter of $3.1\text{\,}{\mathrm{mm}}^{2}$. This facility played a major role in the characterization of one of the large-area silicon detectors used in the Nab experiment, meeting the requirements of segment-by-segment calibration and study.